# The art of modeling nuclear reactions with weakly bound nuclei: status and perspectives

### v1

@@ 7.1 The semiclassical formalism of Alder and Winther
which, when inserted in (118), gives rise to a set of coupled equations for the expansion coefficients:
@@
and the differential cross section
@@
When the couplings are weak, one may solve (120) perturbatively, assuming that $c_{0}\approx 1$ and $c_{n}\ll 1$ for $n>0$. This gives the first-order solution
@@
which is valid only for angles smaller than the grazing one ($\theta<\theta_{\text{gr}}$) and where $a_{0}$ is half of the distance of closest approach in a classical head-on collision, $\xi_{0\rightarrow n}=\frac{(\varepsilon_{n}-\varepsilon_{0})}{\hbar}\frac{a_{0}}{v}$ is the adiabaticity parameter and $f_{\lambda}(\theta,\xi)$ is an analytic function, depending on the kinematical conditions, but independent of the structure of the projectile.
@@
where ${dN_{E\lambda}(E_{x}\theta)}/{d\Omega}$ is the number of equivalent photons per solid angle.
Figure 44, quoted from Ber09 , shows the number of equivalent virtual photons, integrated for impact parameters beyond $b=12.3$ fm, for the collision of a projectile with a $Z=82$ target, for three different incident energies. It can be seen that increasing the incident energy decreases the population of low-lying states and enhances the population of high-energy ones.
@@
7.2 Application to the extraction the electric transition probabilities from Coulomb dissociation data
Coulomb excitation has been used for many years as a powerful spectroscopic tool to extract the electromagnetic response of nuclei to a Coulomb field, provided by a heavy nucleus. For stable nuclei and low bombarding energies (tens of MeV per nucleon), Coulomb excitation is dominated by transitions to low-lying excited states connected by quadrupole and octupole multipolarities with the ground state, whereas at sufficiently high energies, Coulomb excitation is dominated by the so-called giant dipole resonance (GDR), characterized by collective motion of protons and neutrons moving out-of-phase motion. In the case of weakly bound nuclei and, more specifically, for neutron halo nuclei, the differential force of the Coulomb field acting on the valence neutron(s) with respect to the core produces a strong dipole force that may eventually break the system. This mode has been customarily termed as “soft dipole E1” excitation, as opposed to the giant dipole excitation, see Fig. 45.
Figure 45: B(E1) distribution for nucleus 11Li. At low energies, the peak associated to the weakly-bound neutrons forms the so-called “pygmy” or “soft dipole” resonance. Figure taken from Ref. Bro19 .
The concentration of dipole strength at low-excitation energies is a general property of weakly-bound nuclei, particularly those with a neutron-halo structure. This can be verified considering a simple situation of a two-body system described with a single-particle model in which the $B(E1)$ can be readily evaluated numerically. Considering such model, Fig. 46 displays the $B(E1)$ transition for a hypothetical (and unrealistic) system with mass $A=8$ and a nucleon binding energy $E_{b}$ = 0.7 MeV for different assumptions of the initial state orbital angular momentum. In the valence proton case, the strength distribution is shifted to higher energies, together with a damping of the absolute value. The concentration of strength becomes very large near the breakup threshold for an initial state with $\ell=0$.
The aim of a Coulomb dissociation experiment with drip-line nuclei is the extraction of the electric response to the continuum that, in the case of halo nuclei, is expected to be largely dominated by its $E1$ distribution.
An appealing feature of the semiclassical formulas outlined in the previous section is that they provide a neat separation between the structure of the projectile (through $B(E\lambda)$) and the reaction dynamics. This separation allows in principle to extract the ${dB(E\lambda)}/{d\varepsilon}$ distribution from the analysis of small-angle Coulomb dissociation data. For that, one can integrate the double differential cross section (124) up to a maximum angle ($\theta_{\mathrm{max}}$) such that for $\theta<\theta_{\mathrm{max}}$ the breakup can be assumed to be purely Coulomb and hence (124) is valid ($\theta_{\mathrm{max}}<\theta_{\text{gr}}$). This gives:
[…]
Restricting the analysis to the lower cut ($\theta_{\mathrm{c.m.}}<1.3^{\circ}$), and assuming that for these angles the dissociation probability is dominated by the Coulomb dipole couplings, one can use the approximation (127) and extract the $B(E1$) distribution. The result of such an analysis, performed in Fuk04 , is shown in Fig. 48, which shows the $B(E1)$ extracted in Fuk04 (black squares).
In the particular case of 11Be the ground state is known to contain admixtures of $s$ and $d$-waves for the neutron-core relative motion (see Sec. 3.3). As noted above, the $B(E1)$ is largely dominated by the asymptotic part of the ground-state wavefunction and hence most of the contribution is due to the $s$-wave component. Consequently, the magnitude of the $B(E1)$ is roughly proportional to the weight of this component or, more correctly, by the spectroscopic factor associated with the overlap $\langle^{11}\mathrm{Be}(g.s.)|^{10}\mathrm{Be}(0^{+})\otimes 2s_{1/2}\rangle$. This way, Coulomb dissociation experiments constitute complementary spectroscopic tools to transfer or knockout reactions. As in those cases, a possible caveat of this methodology is the dependence of the spectroscopic factor on the assumed single-particle model. A detailed discussion of this technique can be found in recent topical reviews Aum13 ; Nak17 .
If the approximations implied by the EPM are well fulfilled, the extracted $B(E\lambda$) distributions should be independent of the incident energy and the target. This is not always the case, as can be seen by comparing the $B(E1)$ discussed above with that obtained from the analysis of another 11Be+208Pb experiment performed at GSI by Palit et al. Pal03 at the much higher energy of 520 MeV/u, shown in Fig. 48 by the yellow circles. The origin of this discrepancy is further discussed in Sec. 7.6).
@@ 7.3 Dynamic Coulomb polarization potential from the AW theory
Feshbach theory of nuclear reactions provides formal expressions for the contribution on the optical potential of nonelastic channels. These contributions are such that, when added to the bare interaction between the colliding nuclei (that is, the static part of the interaction due to the ground-state densities for the projectile and target) yield the “correct” elastic scattering cross section. In general, this dynamical contribution of the optical potential, termed polarization potential, is nonlocal, and energy and angular momentum dependent. Under suitable approximations, it is possible to derive simple, analytical expressions for specific parts of the polarization potential.
Figure 50: Second-order process describing the excitation and de-excitation of a nucleus moving along a classical trajectory determined by the potential $V_{0}(t)$, due to the action of the potential $V_{1}(t)$.
In the case of Coulomb excitation to either bound or unbound states (breakup) the semiclassical theory of Alder and Winther outlined in the previous section provides a very convenient framework for obtaining these simple forms of the Coulomb polarization potential. One such example is provided by the works of Andrés, Gómez and Nagarajan And94 ; And95 .
These works consider the second-order process in which the projectile is excited to a continuum state due to the dipole Coulomb interaction ($V_{1}(t)$) and then de-excites, returning to the ground state. The situation is schematically depicted in Fig. 50. The amplitude probability for this process is given by:
@@
where $|z\rangle$ denotes the intermediate states populated during the reaction. Then, one requires that this second-order amplitude coincides with the first-order amplitude associated with the polarization potential for all classical trajectories corresponding to a given scattering energy. This gives rise to the following expresssion:
@@
and $\xi$ is the Coulomb adiabaticity parameter corresponding to the excitation to the continuum energy $\varepsilon$ of the nucleus. An important feature of this potential is that when the breakup energy $\varepsilon_{b}$ is large enough, the purely real adiabatic dipole potential given by Eq. (2). In the opposite limit, for small breakup energies, $f\left(\frac{R}{a_{o}}-1,\xi\right)\rightarrow 1$ and $g\left(\frac{R}{a_{o}}-1,\xi\right)\rightarrow 0$, and the polarization potential becomes purely imaginary, depending on $R$ as ${1}/[(R-a_{o})^{2}R].$
@@
Figure 52: Real and imaginary parts of the bare and CDP potentials for 6He+208Pb at $E_{\mathrm{lab}}=22$ MeV computed with dynamic Coulomb polarization potential of Eq. (7.3).
As an example, we show in Fig. 51 the elastic scattering data for 6He+208Pb at 22 MeV compared with the optical model calculations obtained with a bare potential alone (dashed) and with the bare plus the Coulomb dipole polarization (CDP) potential (solid line). For the bare interaction, we have used an optical potential for the nearby projectile 6Li, which has a structure similar to that of 6He but which does not exhibit the strong dipole excitation mechanism. The elastic cross section predicted by the bare potential presents a Fresnel-like diffraction pattern, not observed in the data. The inclusion of the CDP significantly reduces the elastic cross section and suppresses the Fresnel peak. In Fig. 52, the real and imaginary parts of the bare and CDP potentials are shown. Their most remarkable feature is their long range, which was anticipated from the optical model analysis of Sec. 2, and is associated with the long range of the dipole Coulomb couplings, as we have already discussed in the context of the CDCC method.
@@ 7.4 Semiclassical transfer-to-the-continuum model
The semiclassical STC model Bon88 developed by Bonaccorso and Brink is a generalization of the model of Brink and collaborators Has78 ; Has79 ; Mon85 ; Bon87 to the case of final unbound states. These methods provide a more transparent and often analytical interpretation of the quantum-mechanical counterparts. The method is valid for peripheral reactions based on the core spectator model as in the IAV model discussed in previous sections. To obtain an analytic expression for the transfer amplitude the STC replaces the internal state wavefunctions by their asymptotic forms, given analytically in terms of Hankel functions.
Considering the case of a neutron transfer, the starting point of the STC is the semiclassical transfer to the continuum scattering amplitude, given by Bon88
@@
$\phi_{l_{i}}(\boldsymbol{r}_{Cn})\simeq-C_{i}i^{l}\gamma h^{(1)}_{l_{i}}(i\gamma_{i}r_{Cn})Y_{l_{i},m_{i}}(\Omega_{i}).$ (134)
with $C_{i}$ the asymptotic normalization coefficient of the initial bound state wavefunction. The the final (continuum) state is approximated by:
@@
The first and second terms correspond to the elastic and nonelastic breakup contributions discussed in the context of the IAV model. The $\bar{S}_{j_{f}}$ are nucleon-target S-matrices calculated for each nucleon final energy according to the optical model with an energy-dependent optical potential. The $B_{if}$ are kinematical factors depending on the energies ($\varepsilon_{i}$ and $\varepsilon_{f}$), momenta ($\gamma_{i}$ and $k_{f}$), and orbital angular momenta ($l_{i}$ and $l_{f}$) of the initial and final states of a single neutron particle, respectively, and on the incident energy per particle, $mv^{2}/2$ at the distance of closest approach $d$ , which coincides with the projectile impact parameter in a straight line trajectory. Their explicit form can be found in the original works Bon88 .
@@
For sufficiently high incident energies, the $S$-matrix can be evaluated in the eikonal approximation and one can also use the connection between the angular momentum and impact factor to transform the sum in Eq. (136) by an integral over the impact parameter, i.e, $\Sigma_{j_{f}}\to\int d^{2}\boldsymbol{b}_{n}$.
Figure 53: n-10Be relative energy spectrum, including Coulomb and nuclear breakup for the reaction 11Be+12C $\to$ n+10Be+X at 69 A.MeV. Only the contributions from an s initial state with spectroscopic factor C2S= 0.84 are calculated. The triangles are the total calculated result after convolution with the experimental resolution function. The dots are the experimental points from Fuk04 . Figure taken from Ref. Blan07 , with authorization from Elsevier.
An example of the application of the method is given in Fig. 53. It corresponds to the reaction 11Be+12C $\to$ $n$+10Be+12C. The STC calculations for different angular momenta of the final state are compared with the data from Fukuda et alFuk04 .
@@ 7.5 Dynamical Eikonal approximation
The dynamical eikonal approximation (DEA) developed by the group at the ULB Bay05 ; Gol06 unifies the semiclassical time-dependent and eikonal methods. It allows for the calculation of differential cross sections for elastic scattering and breakup in a quantal way by taking into account interference effects.
As in the CDCC method, the idea of the method is to provide an approximate solution of the three-body Schrodinger equation with the effective Hamiltonian (5). The three-body scattering solution of this equation is expressed in the factorized form:
@@
The DEA equation (140) is mathematically equivalent to a time-dependent Schrodinger equation for a straight-line trajectory. It can therefore be solved using similar numerical techniques as in the time-dependent model (see e.g. Kid94 ; Cap03 ) The DEA equation must be solved for each transverse component $\vec{b}$ of the $a$-$A$ coordinate with the initial boundary condition
@@
Despite its formal analogy, it is important to note that there is no semiclassical approximation involved in the derivation of the method and so the coordinates $Z$ and $\vec{b}$ are quantal variables. This enables us to take into account interferences between trajectories.
@@
7.6 Coulomb dissociation revisited: quantum-mechanical and nuclear effects
In Sec. 7 we presented the standard procedure of extracting electric transition probabilities from Coulomb dissociation data, making use of the first-order semiclassical theory of Alder and Winther (i.e., the equivalent photon method). This procedure faces several difficulties, namely
  1. (i)
    in addition to Coulomb couplings, nuclear couplings are always present to a greater or lesser extent;
  2. (ii)
    even in situations in which the nuclear contribution is small, other multipoles, other than the dipole one, may be present;

### v2

@@ 7.1 The semiclassical formalism of Alder and Winther
which, when inserted in Eq. (128), gives rise to a set of coupled equations for the expansion coefficients:
@@
and the differential cross section by:
@@
When the couplings are weak, one may solve Eq. (130) perturbatively, assuming that $c_{0}\approx 1$ and $c_{n}\ll 1$ for $n>0$. This gives the first-order solution
@@
which is valid only for angles smaller than the grazing [^3] The grazing angle refers to the angle at which the projectile interacts with the surface of the target in such a way that the projectile barely “grazes” the surface of the target, rather than fully penetrating or colliding head-on. one ($\theta<\theta_{\text{gr}}$) and where $a_{0}$ is half the distance of closest approach in a classical head-on collision, $\xi_{0\rightarrow n}=\frac{(\varepsilon_{n}-\varepsilon_{0})}{\hbar}\frac{a_{0}}{v}$ is the adiabaticity parameter and $f_{\lambda}(\theta,\xi)$ is an analytic function, depending on the kinematical conditions, but independent of the structure of the projectile.
@@
where ${dN_{E\lambda}(E_{x}\theta)}/{d\Omega}$ is the number of equivalent photons per solid angle. This is typically referred to as Equivalent Photon Method (EPM).
Figure 48, quoted from Ber09 , shows the number of equivalent virtual photons, integrated for impact parameters beyond $b=12.3$ fm (note that $b=a_{0}\cot{(\theta/2)}$), for the collision of a projectile with a $Z=82$ target, for three different incident energies. It can be seen that increasing the incident energy decreases the population of low-lying states and enhances the population of high-energy ones.
@@ 7.2 Dynamic Coulomb polarization potential from the AW theory
As discussed in Sec. 2, the Feshbach theory of nuclear reactions provides a formal expression for the nucleus-nucleus optical potential. This can be expressed as a sum of a bare potential (that is, the static part of the interaction due to the ground-state densities for the projectile and target) and a polarization potential. Under suitable approximations, it is possible to derive simple, analytical expressions for specific parts of the polarization potential. One example is the adiabatic polarization potential describing the effect of the dipole Coulomb interaction discussed in Sec. 2.1.
This adiabatic expression assumes that the excitation energies are large and, therefore, is not applicable to the Coulomb breakup of very weakly bound nuclei, for which the average excitation energies are typically small (of the order of 1 MeV or less).
Figure 50: Second-order process describing the excitation and de-excitation of a nucleus moving along a classical trajectory determined by potential $V_{0}(t)$, due to the action of potential $V_{1}(t)$.
A non-adiabatic (also called dynamic) polarization potential can be derived from the Alder and Winther theory And94 ; And95 . These works consider the second-order process in which the projectile is excited to a continuum state due to the dipole Coulomb interaction ($V_{1}(t)$) and then de-excites, returning to the ground state. The situation is schematically depicted in Fig. 50. The amplitude probability for this process is given by:
@@
where $|z\rangle$ denotes the intermediate states populated during the reaction. Then, one requires that this second-order amplitude coincides with the first-order amplitude associated with the polarization potential for all classical trajectories corresponding to a given scattering energy. This gives rise to the following expression:
@@
where $K_{2i\xi}(2\xi z)$ is a modified Bessel functions of imaginary order and $\xi$ is the Coulomb adiabaticity parameter corresponding to the excitation to the continuum energy $\varepsilon$ of the nucleus. An important feature of this potential is that when the breakup energy $\varepsilon_{b}$ is large enough, the purely real adiabatic dipole potential given by Eq. (4) is recovered. In the opposite limit, for small breakup energies, $f\left(\frac{R}{a_{o}}-1,\xi\right)\rightarrow 1$ and $g\left(\frac{R}{a_{o}}-1,\xi\right)\rightarrow 0$, and the polarization potential becomes purely imaginary, depending on $R$ as ${1}/[(R-a_{o})^{2}R].$
@@
Figure 52: Real and imaginary parts of the bare and CDP potentials for 6He+208Pb at $E_{\mathrm{lab}}=22$ MeV computed with the dynamic Coulomb polarization potential of Eq. (7.2).
As an example, we show in Fig. 51 the elastic scattering data for 6He+208Pb at 22 MeV compared with the optical model calculations obtained with a bare potential alone (dashed) and with the bare plus the Coulomb dipole polarization (CDP) potential (solid line). For the bare interaction, we have used an optical potential for the nearby projectile 6Li, which has a structure similar to that of 6He but which does not exhibit the strong dipole excitation mechanism. The elastic cross section predicted by the bare potential presents a Fresnel-like diffraction pattern, not observed in the data. The inclusion of the CDP significantly reduces the elastic cross section and suppresses the Fresnel peak. This plot shows that the Coulomb polarization effect (encoded here in the form of a polarization potential), is able to account for part of the effects of the breakup channels on the elastic channel. The remaining disagreement with the data can be explained as due to other couplings not accounted for by this Coulomb polarization potential, namely, nuclear couplings and possible interferences between nuclear and Coulomb. These effects are naturally accommodated within the CDCC formalism discussed in 3.
In Fig. 52, the real and imaginary parts of the bare and CDP potentials are shown. Their most remarkable feature is their long range, which was anticipated from the optical model analysis of Sec. 2, and is associated with the long range of the dipole Coulomb couplings, as we have already discussed in the context of the CDCC method.
@@ 7.3 Semiclassical transfer-to-the-continuum model
The Semiclassical Transfer-to-the-Continuum (STC) model Bon88 developed by Bonaccorso and Brink is a generalization of the model of Brink and collaborators Has78 ; Has79 ; Mon85 ; Bon87 to the case of final unbound states. These methods provide a more transparent and often analytical interpretation of the quantum-mechanical counterparts. The method is valid for peripheral reactions based on the core spectator model as in the IAV model discussed in previous sections. To obtain an analytic expression for the transfer amplitude the STC replaces the internal state wavefunctions by their asymptotic forms, given analytically in terms of Hankel functions.
Considering the case of neutron transfer, the starting point of the STC is the semiclassical transfer to the continuum scattering amplitude, given by Bon88
@@
$\phi_{l_{i}}(\boldsymbol{r}_{Cn})\simeq-C_{i}i^{l}\gamma_{i}h^{(1)}_{l_{i}}(i\gamma_{i}r_{Cn})Y_{l_{i},m_{i}}(\Omega_{i}).$ (142)
with $\gamma_{i}^{2}=-2m\varepsilon_{i}/\hbar^{2}$, $C_{i}$ the asymptotic normalization coefficient of the initial bound state wavefunction. The final (continuum) state is approximated by:
@@
The first and second terms correspond to the elastic and nonelastic breakup contributions discussed in the context of the IAV model. $\bar{S}_{j_{f}}$ are nucleon-target S-matrices calculated for each nucleon final energy according to the optical model with an energy-dependent optical potential. $B_{if}$ are kinematical factors depending on the energies ($\varepsilon_{i}$ and $\varepsilon_{f}$), momenta ($\gamma_{i}$ and $k_{f}$), and orbital angular momenta ($l_{i}$ and $l_{f}$) of the initial and final states of a single neutron particle, respectively, and on the incident energy per particle, $mv^{2}/2$ at the distance of closest approach $d$, which coincides with the projectile impact parameter in a straight line trajectory. Their explicit form can be found in the original works Bon88 .
@@
For sufficiently high incident energies, the $S$-matrix can be evaluated in the eikonal approximation and one can also use the connection between angular momentum and impact factor to transform the sum in Eq. (144) to an integral over the impact parameter, i.e, $\Sigma_{j_{f}}\to\int d^{2}\boldsymbol{b}_{n}$.
Figure 53: Experimental and calculated cross section momentum distribution for the proton and neutron knockout reactions of 14O on 9Be at 55 MeV/a. The solid lines are the results obtained by the IAV model and the dashed lines are calculated by the semiclassical transfer to the continuum (TC) method. Adapted from Ref. Jin21 . Experimental data are from Ref. Fla12 .
An example of the application of the method is given in Fig. 53. It corresponds to the knockout reactions 9Be(14O,13O)X 9Be(14O,13N)X at 55 MeV/u. The semiclassical transfer to the continuum calculations are compared to those obtained with the IAV model discussed in Sec. 5.1 and with the experimental data from Fla12 . A nice agreement between the purely quantum mechanical and semiclassical methods is observed.
@@ 7.4 Dynamical Eikonal approximation
The dynamical eikonal approximation (DEA) Bay05 ; Gol06 unifies the semiclassical time-dependent and eikonal methods. It allows for the calculation of differential cross sections for elastic scattering and breakup in a quantal way by taking into account interference effects.
As in the CDCC method, the idea of the method is to provide an approximate solution of the three-body Schrodinger equation with the effective Hamiltonian in Eq. (6). The three-body scattering solution for this equation is expressed in the factorized form:
@@
The DEA equation (148) is mathematically equivalent to a time-dependent Schrodinger equation for a straight-line trajectory. It can therefore be solved using similar numerical techniques as in the time-dependent model (see e.g. Kid94 ; Cap03 ). The DEA equation must be solved for each transverse component $\vec{b}$ of the $a$-$A$ coordinate with the initial boundary condition
@@
Despite its formal analogy, it is important to note that there is no semiclassical approximation involved in the derivation of the method and so the coordinates $Z$ and $\vec{b}$ are quantal variables. This enables to take into account interferences between trajectories.
@@
8 Extraction the electric transition probabilities from Coulomb dissociation data
Coulomb excitation has been used for many years as a powerful spectroscopic tool to extract the electromagnetic response of nuclei to a Coulomb field, provided by a heavy nucleus. For stable nuclei and low bombarding energies (tens of MeV per nucleon), Coulomb excitation is dominated by transitions to low-lying excited states connected by quadrupole and octupole multipolarities to the ground state, whereas at sufficiently high energies, Coulomb excitation is dominated by the so-called giant dipole resonance, characterized by the collective motion of protons and neutrons moving in an out-of-phase motion. In the case of weakly bound nuclei and, more specifically, for neutron halo nuclei, the differential force of the Coulomb field acting on the valence neutron(s) with respect to the core produces a strong dipole force that may eventually break the system. This mode has been customarily termed as “soft dipole E1” excitation, as opposed to the giant dipole excitation, see Fig. 55.
Figure 55: B(E1) distribution for nucleus 11Li. At low energies, the peak associated to the weakly-bound neutrons forms the so-called “pygmy” or “soft dipole” mode. Figure taken from Ref. Bro19 .
The concentration of dipole strength at low-excitation energies is a general property of weakly-bound nuclei, particularly those with a neutron-halo structure. This can be verified considering a simple situation of a two-body system described with a single-particle model in which the $B(E1)$ can be readily evaluated numerically. Considering such a model, Fig. 56 displays the $B(E1)$ transition for a hypothetical (and unrealistic) system with mass $A=8$ and a nucleon binding energy $E_{b}$ = 0.7 MeV for different assumptions of the initial state orbital angular momentum. In the valence-proton case, the strength distribution is shifted to higher energies, together with a damping of the absolute value. The concentration of strength becomes very large near the breakup threshold for an initial state with $\ell=0$. It is worth noting that the presence of this near-threshold peak does not necessarily imply the existence of a resonant state at this energy.
The aim of a Coulomb dissociation experiment with dripline nuclei is to extract the electric response to the continuum, which, in the case of halo nuclei, is expected to be largely dominated by its $E1$ distribution.
8.1 Semiclassical analysis of Coulomb dissociation data
An attractive feature of the semiclassical formulas discussed in Sec. 7 is that they provide a neat separation between the structure of the projectile (through $B(E\lambda)$) and the reaction dynamics. In practical situations, nuclear couplings will be also present and there will not be such a simple connection between the measure breakup cross sections and the underlying ${dB(E\lambda)}/{d\varepsilon}$ distribution.